# Scalable approximation of the transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

### v3

@@ 4.1 Per-iteration Time Complexity
  • EM Algorithm: The E-step computes the latent allocations $z_{ijk}^{(t)}$ for all $i\in n$, $j\in D_{p}$, and $k\in D_{r}$. For each triplet $(i,j,k)$, the E-step costs $\mathcal{O}(nD_{p}^{2}D_{r})$. The M-step performs normalization in $\mathcal{O}(nD_{p}D_{r})$ time. Thus, the complexity of the $t$-th iteration is
    $T_{\text{EM}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r}).$
  • CIRLS Algorithm: After vectorization, the main computational steps are: a) computing the linear predictor $\bm{\mu}^{(t)}=\tilde{\mathbf{X}}\tilde{\bm{\beta}}^{(t)}$ in $\mathcal{O}(nD_{p}D_{r})$ time; b) computing the weight matrix $\mathbf{W}^{(t)}$ in $\mathcal{O}(nD_{r})$ time; c) forming $\mathbf{D}=\tilde{\mathbf{X}}^{\top}\mathbf{W}^{(t)}\tilde{\mathbf{X}}$, which is block diagonal with $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(nD_{p}^{2}D_{r})$ operations; and d) solving the QP problem with $D_{rp}=D_{r}D_{p}$ variables and $\mathcal{O}(D_{rp})$ constraints. Using interior-point methods, the QP solve costs $\mathcal{O}(D_{rp}^{3})=\mathcal{O}(D_{p}^{3}D_{r}^{3})$. The complexity of the $t$-th iteration is
    $T_{\text{CIRLS}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}).$
@@
If EM requires $K_{\text{EM}}$ iterations and CIRLS requires $K_{\text{CIRLS}}$ iterations to converge, the total complexities are
$T_{\text{EM}}$ $=\mathcal{O}(K_{\text{EM}}\cdot nD_{p}^{2}D_{r}),$
$T_{\text{CIRLS}}$ $=\mathcal{O}(K_{\text{CIRLS}}\cdot[nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}]).$
@@ 4.2 Memory Complexity
  • EM Algorithm: The main memory requirements are: a) the data matrices $\mathbf{X}$ and $\mathbf{Y}$, requiring $\mathcal{O}(n(D_{p}+D_{r}))$ space; b) the coefficient matrix $\mathbf{B}$, requiring $\mathcal{O}(D_{p}D_{r})$ space; and c) the latent allocation array $z_{ijk}$, requiring $\mathcal{O}(nD_{p}D_{r})$ space. Thus,
    $M_{\text{EM}}=\mathcal{O}(nD_{p}D_{r}).$
  • CIRLS Algorithm: In addition to the data and coefficients, CIRLS requires: a) the block-diagonal matrix $\mathbf{D}$, which can be stored efficiently as $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(D_{p}^{2}D_{r})$ space; b) the QP solver workspace, typically requiring $\mathcal{O}(D_{p}^{2}D_{r}^{2})$ space for storing constraint matrices and factorizations. Thus,
    $M_{\text{CIRLS}}=\mathcal{O}(nD_{p}D_{r}+D_{p}^{2}D_{r}^{2}).$
@@ 5 Simulation studies
Simulation studies [^4] The experiments were conducted utilising a Dell laptop equipped with Intel Core i5-1053G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. were conducted to compare the computational cost (running time) of the EM and the CIRLS estimation procedures for the regression coefficients of the TFLR. For all cases, both simplicial responses and predictors were generated from a Dirichlet distribution. The sample sizes considered were 1,000 up to 10,000 with an increasing step size of 10,000 and from then the step size increased to 5,000 up to 50,000. For all cases, the tolerance value to check convergence was $\epsilon=10^{-6}$, and the results were averaged over $100$ replicates for each combination of dimensionality and sample size. Two scenarios were considered: a) the responses and the predictor variables are independent and b) they are linearly dependent, with the following configurations:
@@ 5.1 Speed-up factors
Figure 2 visualizes the speed-up factors when the compositions are dependent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 4 up to 227, whereas compared to EM-2 they range from 2 up to 96. When $p=10$, the EM-1 is 9 up 347 times slower, whereas EM-2 is 12-264 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 46-114 times slower than CIRLS when $p=15$, and 40-157 times slower than CIRLS when $p=20$. Notably, there was a discrepancy in the produced KLD values. CIRLS always produced slightly higher KLD than the EM algorithm, the discrepancy was $\approx 10^{-6}-10^{-4}$.
@@ 5.2 Scalability of the EM and CIRLS algorithms
Further, both sets of simulation studies revealed that the EM-2 is slower than EM-1. Using the CIRLS produced coefficients as initial values in the EM-2 algorithm improved its speed, but not significantly.
@@ 6 Conclusions
We adopted the IRLS methodology used in generalised linear models to the case of the TFLR model for simplicial-simplicial regression. For this we exploited the fact that TFLR may be seen as a logistic regression model with simplicial constraints on the regression coefficients. Regardless of the dimensionalities in both the responses and the predictors, this resulted in significant time improvements ranging from 1.3 up to more than 300 times. The estimated scalability rates were linear or sublinear for both algorithms. The remarkable speed-up factors of CIRLS though are accompanied by a small trade-off. When the compositions are dependent, the CIRLS fails by a small amount ($\approx 10^{-6}-10^{-4}$) to reach the minimum KLD value. This discrepancy is negligible for most practical applications, especially for permutation and bootstrap hypothesis testing.
The CIRLS creates opportunities for additional methodological developments in simplicial-simplicial regression. fiksel2022 proposed to minimize the KLD from the observed to the fitted compositions, and tsagris2025 the objective function with the $L_{2}$ and $L_{1}$ norms. $\phi$-divergence regression is another direction, that includes the symmetric KLD, and the Jensen-Shannon divergence (a symmetrized version of the KLD), among others. The CIRLS algorithm could be used to approximate the exact solutions of this type of regression, yielding fast, but approximate solutions, that could also be used as initial values to an exact algorithm.

### v4

@@ 4.1 Per-iteration Time Complexity
  • EM Algorithm: The E-step computes the latent allocations $z_{ijk}^{(t)}$ for all $i\in n$, $j\in D_{p}$, and $k\in D_{r}$. For each triplet $(i,j,k)$, the E-step costs $\mathcal{O}(nD_{p}^{2}D_{r})$. The M-step performs normalization in $\mathcal{O}(nD_{p}D_{r})$ time. Thus, the complexity of the $t$-th iteration is $T_{\text{EM}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r})$.
  • CIRLS Algorithm: After vectorization, the main computational steps are: a) computing the linear predictor $\bm{\mu}^{(t)}=\tilde{\mathbf{X}}\tilde{\bm{\beta}}^{(t)}$ in $\mathcal{O}(nD_{p}D_{r})$ time; b) computing the weight matrix $\mathbf{W}^{(t)}$ in $\mathcal{O}(nD_{r})$ time; c) forming $\mathbf{D}=\tilde{\mathbf{X}}^{\top}\mathbf{W}^{(t)}\tilde{\mathbf{X}}$, which is block diagonal with $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(nD_{p}^{2}D_{r})$ operations; and d) solving the QP problem with $D_{rp}$ variables and $\mathcal{O}(D_{rp})$ constraints. Using interior-point methods, the QP solve costs $\mathcal{O}(D_{rp}^{3})=\mathcal{O}(D_{p}^{3}D_{r}^{3})$. The complexity of the $t$-th iteration is $T_{\text{CIRLS}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3})$.
@@
If EM requires $K_{\text{EM}}$ iterations and CIRLS requires $K_{\text{CIRLS}}$ iterations to converge, the total complexities are $T_{\text{EM}}=\mathcal{O}(K_{\text{EM}}\cdot nD_{p}^{2}D_{r})$ and $T_{\text{CIRLS}}=\mathcal{O}(K_{\text{CIRLS}}\cdot[nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}])$.
@@ 4.2 Memory Complexity
  • EM Algorithm: The main memory requirements are: a) the data matrices $\mathbf{X}$ and $\mathbf{Y}$, requiring $\mathcal{O}(n(D_{p}+D_{r}))$ space; b) the coefficient matrix $\mathbf{B}$, requiring $\mathcal{O}(D_{p}D_{r})$ space; and c) the latent allocation array $z_{ijk}$, requiring $\mathcal{O}(nD_{p}D_{r})$ space. Thus, $M_{\text{EM}}=\mathcal{O}(nD_{p}D_{r})$.
  • CIRLS Algorithm: In addition to the data and coefficients, CIRLS requires: a) the block-diagonal matrix $\mathbf{D}$, which can be stored efficiently as $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(D_{p}^{2}D_{r})$ space; b) the QP solver workspace, typically requiring $\mathcal{O}(D_{p}^{2}D_{r}^{2})$ space for storing constraint matrices and factorizations. Thus, $M_{\text{CIRLS}}=\mathcal{O}(nD_{p}D_{r}+D_{p}^{2}D_{r}^{2})$.
@@ 5 Simulation studies
Simulation studies [^4] The experiments were conducted utilizing a Dell laptop equipped with Intel Core i5-1035G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. were conducted to compare the computational cost (running time) of the EM and the CIRLS estimation procedures for the regression coefficients of the TFLR. For all cases, both simplicial responses and predictors were generated from a Dirichlet distribution. The sample sizes considered were 1,000 up to 10,000 with an increasing step size of 1,000 and from then the step size increased to 5,000 up to 50,000. For all cases, the tolerance value to check convergence was $\epsilon=10^{-6}$, and the results were averaged over $100$ replicates for each combination of dimensionality and sample size. Two scenarios were considered: a) the responses and the predictor variables are independent and b) they are linearly dependent, with the following configurations:
@@ 5.1 Speed-up factors
Figure 2 visualizes the speed-up factors when the compositions are dependent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 4 up to 84, whereas compared to EM-2 they range from 8 up to 326. When $p=10$, the EM-1 is 15 up 257 times slower, whereas EM-2 is 27-445 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 40-165 times slower than CIRLS when $p=15$, and 22-155 times slower than CIRLS when $p=20$. Notably, there was a discrepancy in the produced KLD values. CIRLS always produced slightly higher KLD than the EM algorithm, the discrepancy was $\approx 10^{-5}-10^{-4}$.
@@ 5.2 Scalability of the EM and CIRLS algorithms
5.3 Theoretical justification for KLD discrepancy
The small discrepancy between EM and CIRLS arises from CIRLS’s use of local quadratic approximations via Taylor expansion. While EM maximizes the exact KLD objective, CIRLS solves a sequence of quadratic programs that approximate the KLD surface around the current iterate. The negligible magnitude of this discrepancy ($\approx 10^{-5}-10^{-4}$) demonstrates that the second-order Taylor approximation is highly accurate in the TFLR setting, making CIRLS a valid and efficient alternative to EM. This is a well-understood trade-off in optimization: Newton-type methods (like IRLS/CIRLS) converge faster but use approximations, while EM guarantees exact monotonic improvement but requires more iterations.
A critical observation is that the Hessian matrix $\mathbf{H}$ is singular due to the simplex constraints. Specifically, for each row of $\bm{B}$, we have the constraint $\sum_{k=1}^{D_{r}}B_{jk}=1$, which means the parameters lie in a $(D_{r}-1)$-dimensional subspace. This introduces one zero eigenvalue per row, making the full Hessian rank-deficient. The rank of the Hessian is $\text{rank}(\bm{H})=D_{p}\cdot(D_{r}-1)<D_{p}\cdot D_{r}$. This means $\bm{H}$ has exactly $D_{p}$ zero eigenvalues (one per simplex constraint).
Both EM and CIRLS solve a constrained optimization problem, for which the Karush-Kuhn-Tucker (KKT) conditions (karush1939minima, gale1951linear) at optimality are
$\nabla\ell(\bm{B}^{*})+\sum_{j=1}^{D_{p}}\lambda_{j}\nabla g_{j}(\bm{B}^{*})+\bm{\nu}=\bm{0},$
where:
  • $g_{j}(\bm{B})=\sum_{k=1}^{D_{r}}B_{jk}-1=0$ is the $j$-th equality constraint (the simplex constraint for the $j$-th row).
  • $\lambda_{j}\in\bm{R}$ is the Lagrange multiplier for the $j$-th equality constraint.
  • $\nabla g_{j}(\bm{B})$ is the gradient of the $j$-th constraint; when $\bm{B}$ is vectorized, this is a vector with 1’s in positions corresponding to $B_{j1},\ldots,B_{jD_{r}}$ and 0’s elsewhere.
  • $\bm{\nu}\geq\bm{0}$ is the vector of Lagrange multipliers for the inequality constraints $B_{jk}\geq 0$.
The proper framework is to consider the projected Hessian $\bm{H}_{\text{proj}}$ onto the tangent space of the constraints. For the simplex constraints, this is the Hessian restricted to the $(D_{r}-1)$-dimensional subspace for each component. Let $\bm{P}$ be the projection matrix onto the constraint manifold. Then:
$\bm{H}_{\text{proj}}=\bm{P}^{T}\bm{H}\bm{P}.$
For problems involving simplicial data, $\bm{H}_{\text{proj}}$ is typically positive definite with minimum eigenvalue $\lambda_{\min}^{\text{proj}}>0$ in the constraint subspace. For two feasible points (matrices) $\bm{B}_{1}$ and $\bm{B}_{2}$ (both satisfying the simplex constraints), the strong convexity in the constraint subspace gives
$\ell(\bm{B}_{1})-\ell(\bm{B}_{2})\geq\nabla\ell(\bm{B}_{2})^{\top}(\bm{B}_{1}-\bm{B}_{2})+\frac{\lambda_{\min}^{\text{proj}}}{2}\|\bm{P}(\bm{B}_{1}-\bm{B}_{2})\|^{2}.$
Since both solutions are feasible, $\bm{P}(\bm{B}_{1}-\bm{B}_{2})=\bm{B}_{1}-\bm{B}_{2}$ (the difference vector is already in the constraint subspace). Let $\Delta\ell=\ell(\bm{B}^{*}_{\text{EM}})-\ell(\bm{B}^{*}_{\text{CIRLS}})$ be the observed KLD difference. Then:
$\|\bm{B}^{*}_{\text{EM}}-\bm{B}^{*}_{\text{CIRLS}}\|\leq\sqrt{\frac{2|\Delta\ell|}{\lambda_{\min}^{\text{proj}}}}.$
For the TFLR model, the projected Hessian can be approximated by
$\bm{H}_{\text{proj}}\approx\text{diag}\left(\bm{X}^{\top}\bm{V}_{1}^{\text{proj}}\bm{X},\ldots,\bm{X}^{\top}\mathbf{V}_{D_{r}-1}^{\text{proj}}\bm{X}\right),$
where $\bm{V}_{k}^{\text{proj}}$ are the projected weight matrices, and the minimum eigenvalue is aprproximately given by
$\lambda_{\min}^{\text{proj}}\approx\mathcal{O}\left(n\cdot\min_{i,j,k}\left[\frac{x_{ij}^{2}\cdot y_{ik}(1-y_{ik})}{(\sum_{j}x_{ij}B_{jk})^{2}}\right]\right).$
For well-conditioned simplicial data (no components too close to 0 or 1), this is approximately $\lambda_{\min}^{\text{proj}}\approx\mathcal{O}(n)$. Given the observed KLD differences of $|\Delta\ell|\approx 10^{-5}-10^{-4}$ in the dependent case, and using $\lambda_{\min}^{\text{proj}}\approx c\cdot n$ where $c$ is a problem-dependent constant (typically $c\in[0.1,1]$)
$\|\bm{B}^{*}_{\text{EM}}-\bm{B}^{*}_{\text{CIRLS}}\|_{F}\lesssim\mathcal{O}\left(\sqrt{\frac{10^{-4}}{c\cdot n}}\right)$
For example, for $n=1,000$ and assuming $c\approx 0.5$ we obtain
$\|\bm{B}^{*}_{\text{EM}}-\bm{B}^{*}_{\text{CIRLS}}\|_{F}\lesssim\mathcal{O}\left(\sqrt{\frac{10^{-4}}{500}}\right)\approx 4.5\times 10^{-4}$
5.4 Permutation testing
We examined the effect of this discrepancy on the permutation testing. Using the same simulation studies described earlier, we performed permutation testinf for independence between the simplicial responses and the simplicial predictors. We performed 100 iterations and for each iteration we computed the p-value based on the EM algorithm and based on the CIRLS algorithm using 300 permutations. For each iteration, we computed the correlation coefficient between the permuted KLD values of the two algorithms. The correlation coefficient was always equal to 1 and the p-values were always the same.
@@ 6 Conclusions
We adapted the IRLS methodology employed in generalized linear models to the TFLR framework for simplicial-simplicial regression. This adaptation exploits the formal equivalence between TFLR and logistic regression with simplicial constraints imposed on the coefficient matrix. Irrespective of the dimensionalities of both response and predictor simplices, this approach yields substantial computational improvements, with speed-up factors ranging from $6\times$ to over $300\times$. Both algorithms exhibit linear or sublinear scalability with respect to sample size and dimensionality. However, the exceptional computational efficiency of CIRLS entails a modest trade-off: when compositions exhibit dependence structure, CIRLS attains KLD values that deviate marginally ($\approx 10^{-6}$$10^{-4}$) from the global minimum.
This discrepancy is negligible for most practical applications, particularly for permutation-based and bootstrap hypothesis testing procedures. Our simulation studies demonstrate that permutation-based $p$-values derived from EM and CIRLS exhibit strong concordance.
The reduced computational burden of CIRLS confers substantial advantages in these contexts, especially when analyzing large-scale data (with respect to sample size, dimensionality, or both). These benefits extend to scenarios involving multiple datasets or extensive simulation studies, where computational efficiency becomes paramount. Moreover, CIRLS is naturally amenable to parallelization, and through judicious programming strategies (e.g., pre-computation of invariant quantities), permutation and bootstrap testing procedures can achieve exceptional computational efficiency.
The CIRLS framework enables several promising methodological extensions in simplicial-simplicial regression. fiksel2022 proposed minimizing the KLD from observed to fitted compositions, while tsagris2025 explored objective functions based on $L_{2}$ and $L_{1}$ norms. $\phi$-divergence regression represents another natural extension, encompassing the symmetric KLD and Jensen-Shannon divergence (a symmetrized variant of the KLD), among others. The CIRLS algorithm can provide approximate solutions to such regression problems, yielding computationally efficient initial values for subsequent refinement via the EM algorithm.
Declarations
Conflict of interest statement: The author declares no conflict of interest.
Funding Information: This research received no funding.
Data availability: This research has used no data.